# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ 5.1. Data-Centric Supervised Finetuning
Following these steps, supervised fine-tuning can significantly enhance the model’s performance in specific domains and tasks.
@@ 5.1.1. SFT Data Collection and Processing
The sources of supervised fine-tuning (SFT) datasets are similar to those of pre-training datasets, consisting of X-text pairs, as introduced in Appendix A.1. We list commonly used SFT models and their data constitution in Table 1, Table 2, and Table 3 for image, video, and audio modalities, respectively.
@@ 5.1.2. Instruction-Response Pairs Generation
Question answer task requires the model to answer questions based on the given information on the X modality. This kind of task includes simple questions answers, and multiple choices. The Most well-known question-answer task is the visual question answer (VQA) task for visual modality. For VQA tasks, previous research on computer vision has developed a large amount of VQA datasets, and recent work on MLLMs leverages these VQA datasets for constructing instruction-response datasets. For example, LLaVA-1.5 (Liu et al., 2023c) leverage GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), OKVQA (Marino et al., 2019) and VQAv2 (Goyal et al., 2019) to enrich their instruction response datasets. They format these datasets in a uniform template as ”Answer the question using a single word or phrase.” TextVQA (Singh et al., 2019) is another source used for constructing instruction-response datasets which contain text information in images. M3IT (Li et al., 2023l) utilize VQA datasets, such as VQA-v2 (Goyal et al., 2019), Shapes VQA (Andreas et al., 2016) and DocVQA (Mathew et al., 2021), to develop a multimodal instruction tuning dataset that contains 2.4 million instances. In addition to visual question-answer tasks, instruction tuning to improve model performance in the video question answer task also leverages previous video question answer datasets, such as MSRVTT-QA (Xu et al., 2017), iVQA (Yang et al., 2021), MSVD-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019).
@@
Shikra-RD (Chen et al., 2023j) focuses on enhancing the models’ ability for referential dialogue problems, i.e., users point to specific areas and ask questions. They collect public VQA datasets, image captioning data, and several datasets with existing positional annotations to build instruction tuning datasets. Leveraging captions from the Flickr30K (Young et al., 2014) image-caption dataset and using GPT-4, they obtain high-quality RD annotations. PVIT (Chen et al., 2023h) converts the traditional VQA datasets such as GQA and OCR into instruction-response datasets with region-level information. MiniGPT-V2 (Chen et al., 2023m) designs a multi-task instruction template, following the conversation template from LLaMA-2 (Li et al., 2023g). They add task identifier tokens and spatial location representations to identify the task and spatial information in the image. They build five types of instruction-response datasets, including LLaVA-Instruct-150K, grounded image captions from Flickr30K (Young et al., 2014) with direct instruction and object parsing, multi-round conversations, and text-only unnatural instruction datasets (Honovich et al., 2022b). CogVLM (Wang et al., 2024c) provides an IT dataset called CogVLM-SFT-311K. To construct this dataset, they manually select high-quality IT data from MiniGPT-4 and integrate it with LLaVA-Instruct-150K, translating it into Chinese. They then manually correct any noise in their collected instruction tuning datasets and translate them back into English. In their model training, they also leverage VQA datasets and bounding box instruction tuning datasets generated from four types of datasets: grounded captioning (GC) datasets with box-noun phrase pairs in each image, referring expression generation (REG) datasets with box and explanation, referring expression comprehension (REC) datasets with boxes annotated for text, and grounded visual question answering (GroundedVQA) datasets. LLaVA1.5 (Liu et al., 2023c) leverage region level VQA datasets such as Visual Genome (Krishna et al., 2017) and RefCOCO (Kazemzadeh et al., 2014; Mao et al., 2016) in their IT datasets construction.
@@
Another type of downstream task is the classification task, where the data is classified into either a label from a candidate pool or an open-label. M3IT (Li et al., 2023l) leverages various datasets for image classification, including ImageNet (Russakovsky et al., 2015), Grounded Object Identification (COCO-GOI)(Lin et al., 2014), COCO-Text(Veit et al., 2016), Image Text Matching (COCO-ITM)(Lin et al., 2014), e-SNLI-VE(Kayser et al., 2021), Multi-modal Fact Checking (Mocheg)(Yao et al., 2023), and IQA(Duanmu et al., 2021). Additionally, for image classification tasks, InstructBLIP (Dai et al., 2024) uses HatefulMemes (Kiela et al., 2020), a binary classification dataset for detecting hateful content in memes, to generate an instruction-response dataset. X-InstructBLIP (Panagopoulou et al., 2023) further extends this approach to audio classification by using AudioSet (Gemmeke et al., 2017). The instructions used are straightforward sentences asking for the classification of the audio, such as ”Classify the following audio”. And the response is the class of the original data.
@@ 5.1.3. SFT Data Selection
External model-based methods frequently utilize external models to evaluate the quality, diversity, and complexity of data. These models serve various functions, from scoring data to enhancing its characteristics. For instance, Du et al. (2023b) leverage DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023f) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023e) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023c) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
For Multimodal instruction tuning datasets, MLLMs are commonly used to obtain or select high-quality instruction tuning data. LLaVA1.5 (Liu et al., 2023c) is the pioneer in leveraging text-only GPT-4 to expand COCO (Lin et al., 2014) bounding box and caption dataset to a multimodal instruction-following dataset. Following LLaVA, ShareGPT-4v (Chen et al., 2023e) use GPT-4v to improve part of LLaVA’s instruction tuning data.
@@
For self-instruction-based methods, the evaluation and selection of data do not require the involvement of any external models. Li et al. (2023m) proposes a self-instruction method for LLMs to autonomously identify and select challenging data. Li et al. (2023d) identifies high-quality data as those that, when included as part of a few-shot example, can enhance the model’s performance. Kung et al. (2023) proposes a novel task-level uncertainty metric that measures the sensitivity of LLMs to instruction perturbations for a task, to identify high-quality data, and combines this with the concept of active learning to iteratively select data. Liu et al. (2024c) scores data at three granular levels: token, sentence, and model. Each level’s scoring is calculated based on the previous level, with the model-level score serving as the final score.
@@ 5.2. Data-Centric Human Preference Alignment
These prompts encompass multiple tasks such as text generation, question answering, dialogue, summarization, and more. Human taggers are then asked to rank the responses to each prompt, constructing a Q&A dataset used to train the reward model (RM). Subsequently, the fine-tuning stage of the LLM is framed as a reinforcement learning (RL) problem using the trained reward model. In LLaMa2 (Li et al., 2023g), human preferences are further divided into fine-grained aspects, with two separate reward models trained to rank helpfulness and safety, respectively. Further work also aims to improve the models’ generation abilities from these aspects.
@@
To offer a thorough data-centric viewpoint for assessing MLLMs, this section starts with two key aspects. First, we review widely-used methods for evaluating data quality. Then, we provide an overview of MLLMs evaluation datasets, highlighting their collection, processing, and characteristics.
@@ 6.1.1. Dataset Diversity
Dataset diversity is crucial for developing robust and generalizable machine learning models. A diverse dataset enables the model to manage various scenarios and minimizes the risk of bias. Some studies (Heusel et al., 2017; Sajjadi et al., 2018) require a reference distribution or dataset. For instance, Heusel et al. (2017) measures the Wasserstein-2 distance between two Gaussian distributions: one fitted to the embeddings of the reference sample and the other to the embeddings of the sample being evaluated for diversity.
Other papers use similarity scores to define diversity. For example, Shen et al. (2019); Fomicheva et al. (2020) utilize the average pairwise similarity score or its complement, the average dissimilarity. Fomicheva et al. (2020) propose a two-metric evaluation paradigm using precision and recall, with precision measuring quality and recall assessing diversity in terms of coverage of the reference distribution. The Vendi Score (Friedman and Dieng, 2023; Pasarkar and Dieng, 2023; YEH et al., 2023) stands out from existing diversity evaluation metrics by being a reference-free, flexible, and interpretable metric that measures internal diversity without comparing it to a reference distribution. Its reliance on a user-defined similarity function allows for broad applicability across domains, incorporating correlations between features while being computationally efficient and unsupervised. This makes it a valuable tool for diverse machine learning applications.
@@ 6.1.2. Dataset Quality
Dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables models to learn effectively and make precise predictions. Evaluating data quality can be approached from multiple perspectives. TRUE, as presented by Honovich et al. (2022a), provides a comprehensive study evaluating factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, the study demonstrates the robust performance of large-scale NLI and QG-QA methods across various tasks.
Ensuring dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables effective learning and precise predictions. There are several approaches to evaluate data quality. The study TRUE (Honovich et al., 2022a) offers a thorough evaluation of factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, it highlights the robust performance of large-scale NLI and QG-QA methods across different tasks. Object Hallucination (Rohrbach et al., 2018) introduces CHAIR (Caption Hallucination Assessment with Image Relevance), which measures the proportion of words in a generated caption that correspond to objects present in the image, using ground truth sentences and object segmentations. This method helps evaluate the hallucination of the data. FAITH SCORE (Jing et al., 2023) assesses the faithfulness of generated answers from large vision-language models (LVLMs). It identifies descriptive statements, extracts atomic facts, and checks their consistency with input images. This structured process ensures that generated answers align closely with the visual content, providing a comprehensive evaluation of faithfulness in LVLM outputs.
@@ 6.1.3. Dataset Similarity
In machine learning theory, performance improves when the training data distribution closely matches the testing data distribution. Therefore, it is essential to evaluate data distribution distances both theoretically and empirically. Common metrics for measuring distribution similarity include Euclidean distance, KL-divergence, CORAL Loss (Sun and Saenko, 2016), Wasserstein distance, and MMD distance (Jiang et al., 2022). MAUVE Scores (Pillutla et al., 2023) serve as a comparison measure between pairs of distributions, such as those encountered in the generative modeling of text or images. This metric provides statistical bounds and extensive experiments to demonstrate its effectiveness.
@@ 6.2. Evaluation Datasets for MLLMs
In this section, we summarize and compare the commonly used evaluation datasets for MLLMs based on the evaluation tasks.
6.2.1. Captioning Tasks
Captioning datasets act as benchmarks to evaluate a model’s fundamental ability to understand information from different modalities. They can be used to assess the zero-shot or few-shot learning capabilities of MLLMs in multimodality tasks, offering insights into their ability to generalize and adapt to new scenarios with limited training examples.
For image captioning tasks, numerous datasets are available to assess a model’s capability in image understanding. MS-COCO (Chen et al., 2015) is one of the most frequently used datasets for both training and evaluating image captioning. Flickr30K (Young et al., 2014) is also extensively utilized to evaluate models’ capabilities across Zero/Few/Full-shot scenarios. The Karpathy split for MS-COCO and Flickr30K (Karpathy and Fei-Fei, 2015) is commonly adopted for dividing training and testing data in the training and evaluation of MLLMs. Nocaps (Agrawal et al., 2019) was developed to overcome the limitations of MS-COCO by focusing on objects not present in MS-COCO captions, serving as a valuable complement to the MS-COCO dataset.
For video captioning, the MSVD dataset (Chen and Dolan, 2011) is a popular choice for evaluating action recognition and video description tasks (Venugopalan et al., 2015), comprising 2,089 video segments and 85,550 English descriptions, with a standard split provided by Venugopalan et al. (2015). The MSRVTT dataset (Xu et al., 2016) includes 10,000 web video clips totaling 41.2 hours, each annotated with approximately 20 natural sentences. The DiDeMo dataset (Anne Hendricks et al., 2017) contains over 10,000 unedited personal videos, each associated with 3-5 pairs of descriptions pinpointing specific moments. The VATEX dataset (Wang et al., 2019) includes more than 41,250 videos paired with 825,000 captions in both English and Chinese, featuring over 206,000 English-Chinese parallel translation pairs. Additionally, the TVC dataset (Lei et al., 2020) includes both validation and private test components for further evaluation. These datasets are instrumental in evaluating both video captioning and video information retrieval capabilities.
For evaluating audio captioning, the two most commonly used datasets are AudioCaps and Clotho. The AudioCaps dataset (Kim et al., 2019) features 46,000 pairs of audio clips and human-written text descriptions. Meanwhile, the Clotho dataset (Drossos et al., 2020) includes 4,981 audio samples, each lasting between 15 to 30 seconds, accompanied by 24,905 captions that range from 8 to 20 words in length. Both datasets are instrumental in assessing a model’s audio understanding capabilities.
6.2.2. Question Answering Tasks
Evaluation datasets for question answering (QA) are crucial for assessing the performance of MLLMs that handle diverse data types like text, images, and audio. These datasets require models to analyze and combine information from multiple sources to respond to relevant questions accurately.
Visual Question Answering (VQA) evaluation dataset tests the MLLM’s ability to integrate and interpret both text and image data. These datasets contain images paired with questions that require visual understanding to answer. Here are several commonly used VQA evaluation datasets: VQAv2 (Goyal et al., 2019) contains 13 million answers associated with roughly 200,000 images from the COCO dataset. It also contains multi images question answer pairs that can test models’ multi image understanding abilities. GQA (Hudson and Manning, 2019) dataset is designed to address deficiencies in previous visual question answering (VQA) datasets by fostering advanced visual reasoning and reducing biases. It utilizes over 22 million questions generated from 113K real-world images annotated with detailed scene graphs, detailing objects, attributes, and relationships. OKVQA (Marino et al., 2019) was also created to address the limitations of existing VQA benchmarks by focusing on knowledge-based visual question answering. This dataset consists of more than 14,000 questions that require external knowledge to answer, covering various categories such as science & technology, history, and sports. Vizwiz (Gurari et al., 2018) focuses on collecting visual questions from blind users, presenting unique challenges such as image quality and conversational question styles. POPE (Li et al., 2023b) focuses on evaluating object hallucination in vision-language models, using 6,136 binary questions to test if models hallucinate non-existent objects in images from MS-COCO and other sources.
Video question-answering evaluation datasets are developed to assess the capabilities of models in video-to-text translation and overall video comprehension. Key datasets include MSVD-QA (Xu et al., 2017) and MSRVTT-QA (Xu et al., 2017), derived from the MSVD (Chen and Dolan, 2011) and MSRVTT (Xu et al., 2016) video captioning datasets, respectively, which are frequently utilized for video QA tasks. For more intricate web videos, the ActivityNet-QA (Yu et al., 2019) dataset, containing 58,000 QA pairs from 5,800 complex videos, tests models’ abilities to understand complex video content. The TGIF-QA (Jang et al., 2017) dataset, with 165,165 QA pairs from 71,741 animated GIFs, evaluates spatio-temporal reasoning and visual question-answering skills in videos. LSMDC (Rohrbach et al., 2017), which includes 118,114 sentences aligned with clips from 202 movies, is used for assessing video description generation and movie QA abilities. Additionally, MoVQA (Rohrbach et al., 2017), featuring 21,953 manually annotated QA pairs from 100 diverse movies, is aimed at evaluating the understanding of long-form videos over various temporal durations, focusing on complex and extended video content comprehension.
Audio question-answering evaluation datasets are designed to test audio-to-text translation and the ability to understand audio. The ClothoAQA (Lipping et al., 2022) dataset, derived from the Clotho dataset, includes 1,991 audio files, each lasting between 15 to 30 seconds, with six different questions per audio file collected. The MUSIC-AVQA (Li et al., 2022c) dataset contains over 45,000 question-answer pairs extracted from more than 9,000 videos, and over 150 hours of content. It uses 33 question templates across 9 types of questions to support spatio-temporal reasoning in audio-visual scenarios.
6.2.3. Perception and Reasoning Tasks
Perception and reasoning are crucial tests for assessing the general abilities of machine learning language models (MLLMs). TextVQA (Singh et al., 2019) consists of 28,408 images with questions that require reading and reasoning about text in the image. ChartQA (Masry et al., 2022) offers diverse questions based on real-world charts, testing visual and logical reasoning. AI2D (Kembhavi et al., 2016) includes over 5,000 grade school science diagrams with extensive annotations to evaluate diagram interpretation skills. ScienceQA(Lu et al., 2022) presents around 21,000 multimodal questions from science curricula, emphasizing multi-hop reasoning. MathVista (Lu et al., 2023a) comprises 6,141 image-text pairs to benchmark mathematical reasoning in visual contexts. MMVet (Yu et al., 2023) includes 200 images with questions testing large multimodal models across six capabilities. MMBench (Liu et al., 2023a) assesses 20 distinct abilities with 3,000 questions, ranging from object localization to social reasoning. LLaVAW (Liu et al., 2024b) features 24 images with diverse visual content and complex reasoning tests. MME (Fu et al., 2024) focuses on 14 subtasks covering perception and cognition, such as text translation and arithmetic. Lastly, MVBench (Li et al., 2023h) contains 20 video tasks that challenge comprehension of spatial and temporal dynamics, designed to evaluate advanced understanding skills.
There are several datasets designed to evaluate MLLMs’ spatial reasoning abilities. VSR (Liu et al., 2023b) comprises over 10,000 natural text-image pairs with 66 types of spatial relations, making it an excellent resource for evaluating spatial reasoning. The datasets RefCOCO, RefCOCO+, and RefCOCOg (Kazemzadeh et al., 2014; Mao et al., 2016) are crucial for assessing how well models understand and ground natural language expressions within visual contexts, featuring images paired with expressions that describe specific objects or areas. These datasets, which vary in expression collection and annotations, challenge models to integrate vision and language by capturing detailed spatial relationships and contextual nuances. Additionally, the GRIT (Gupta et al., 2022) benchmark addresses seven vision tasks using multiple data sources to enhance referring expression grounding. DocVQA (Mathew et al., 2021) focuses on visual question answering within document images, utilizing a web-based tool to generate its 50,000 questions based on over 12,000 document images. OCR-VQA (Mishra et al., 2019) contains 207,572 images of book covers, with over a million question-answer pairs, testing models’ ability to understand text in document images and perform visual question answering. Lastly, SEED-Bench (Li et al., 2023i) evaluates spatial and temporal comprehension in multimodal contexts, serving as a platform for testing generative comprehension capabilities.
For video reasoning, there exist numerous evaluation datasets. EgoSchema (Mangalam et al., 2024) features over 5,000 very long-form video language understanding questions derived from 250 hours of diverse egocentric video data. VideoChatGPT (Maaz et al., 2023), employed to tackle challenges in video-based conversation models, addresses aspects like temporal understanding, spatial consistency, and contextual comprehension. Both EgoSchema (Mangalam et al., 2024) and VideoChatGPT (Maaz et al., 2023) offer the means to assess models on longer and more intricate video content. Charades-STA (Gao et al., 2017) comprises approximately 10,000 videos, each annotated with temporal activity details across 157 activity categories and multiple video-level descriptions. Additionally, QVHighlights (Lei et al., 2021) contains over 10,000 YouTube videos covering various topics such as everyday activities, travel, and social and political events, facilitating the evaluation of MLLMs’ abilities in detecting moments and generating highlights.
6.2.4. Other Noteworthy Tasks
There are also several noteworthy tasks not mentioned above. One of the traditional tasks is the classification task. For the classification task, InstructBLIP uses HatefulMemes (Kiela et al., 2020), a binary hateful content classification dataset for memes classification. ImageNet-1K (Deng et al., 2009) is also a widely used dataset to test models’ image classification ability. While X-InstructBLIP extends this to ESC50 (Piczak, 2015) for audio classification and ModelNet40 (Wu et al., 2015) for 3D classification. Furthermore, M3IT (Li et al., 2023l) use Something-Something (Goyal et al., 2017) for video action classification. Multimodal dialogue task is different from question answering task because it requires multiple rounds of conversion. TouchStone (Bai et al., 2023b) explores multimodal dialogue through detailed image annotations, addressing the challenge of evaluating LLMs in open-ended dialogues.
@@ 6.2.4. Other Noteworthy Tasks
As elaborated in Section 6.1, despite the introduction of various data evaluation metrics, there remains a notable absence of comprehensive metrics tailored specifically for evaluating multimodal data. The assessment of multimodal data presents heightened challenges owing to its diverse array of data types, often encompassing multiple tasks and modalities (Liu et al., 2023c; Chen et al., 2024c). The quality of such data can be impacted by the individual models corresponding to each modality and their alignments. Hence, it becomes imperative to scrutinize the quality of multimodal data from multiple dimensions.
In addition to relying solely on human-defined features, enhancing data quality can be achieved by leveraging statistical metrics and methodologies, such as scrutinizing distributions. The notion of domain adaptation offers theoretical underpinnings for matching distributions (Jiang et al., 2022). Consequently, assessing distributions emerges as a promising avenue, both in terms of theoretical frameworks and empirical validation.

### v2

@@ 5.1.2. Instruction-Response Pairs Generation
Question answer task requires the model to answer questions based on the given information on the X modality. This kind of task includes simple questions answers, and multiple choices. The Most well-known question-answer task is the visual question answer (VQA) task for visual modality. For VQA tasks, previous research on computer vision has developed a large amount of VQA datasets, and recent work on MLLMs leverages these VQA datasets for constructing instruction-response datasets. For example, LLaVA-1.5 (Liu et al., 2023c) leverage GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), OKVQA (Marino et al., 2019) and VQAv2 (Goyal et al., 2019) to enrich their instruction response datasets. They format these datasets in a uniform template as ”Answer the question using a single word or phrase.” TextVQA (Singh et al., 2019) is another source used for constructing instruction-response datasets which contain text information in images. M3IT (Li et al., 2023k) utilize VQA datasets, such as VQA-v2 (Goyal et al., 2019), Shapes VQA (Andreas et al., 2016) and DocVQA (Mathew et al., 2021), to develop a multimodal instruction tuning dataset that contains 2.4 million instances. In addition to visual question-answer tasks, instruction tuning to improve model performance in the video question answer task also leverages previous video question answer datasets, such as MSRVTT-QA (Xu et al., 2017), iVQA (Yang et al., 2021), MSVD-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019).
@@
Shikra-RD (Chen et al., 2023j) focuses on enhancing the models’ ability for referential dialogue problems, i.e., users point to specific areas and ask questions. They collect public VQA datasets, image captioning data, and several datasets with existing positional annotations to build instruction tuning datasets. Leveraging captions from the Flickr30K (Young et al., 2014) image-caption dataset and using GPT-4, they obtain high-quality RD annotations. PVIT (Chen et al., 2023h) converts the traditional VQA datasets such as GQA and OCR into instruction-response datasets with region-level information. MiniGPT-V2 (Chen et al., 2023m) designs a multi-task instruction template, following the conversation template from LLaMA-2 (Li et al., 2023f). They add task identifier tokens and spatial location representations to identify the task and spatial information in the image. They build five types of instruction-response datasets, including LLaVA-Instruct-150K, grounded image captions from Flickr30K (Young et al., 2014) with direct instruction and object parsing, multi-round conversations, and text-only unnatural instruction datasets (Honovich et al., 2022b). CogVLM (Wang et al., 2024b) provides an IT dataset called CogVLM-SFT-311K. To construct this dataset, they manually select high-quality IT data from MiniGPT-4 and integrate it with LLaVA-Instruct-150K, translating it into Chinese. They then manually correct any noise in their collected instruction tuning datasets and translate them back into English. In their model training, they also leverage VQA datasets and bounding box instruction tuning datasets generated from four types of datasets: grounded captioning (GC) datasets with box-noun phrase pairs in each image, referring expression generation (REG) datasets with box and explanation, referring expression comprehension (REC) datasets with boxes annotated for text, and grounded visual question answering (GroundedVQA) datasets. LLaVA1.5 (Liu et al., 2023c) leverage region level VQA datasets such as Visual Genome (Krishna et al., 2017) and RefCOCO (Kazemzadeh et al., 2014; Mao et al., 2016) in their IT datasets construction.
@@
Another type of downstream task is the classification task, where the data is classified into either a label from a candidate pool or an open-label. M3IT (Li et al., 2023k) leverages various datasets for image classification, including ImageNet (Russakovsky et al., 2015), Grounded Object Identification (COCO-GOI)(Lin et al., 2014), COCO-Text(Veit et al., 2016), Image Text Matching (COCO-ITM)(Lin et al., 2014), e-SNLI-VE(Kayser et al., 2021), Multi-modal Fact Checking (Mocheg)(Yao et al., 2023), and IQA(Duanmu et al., 2021). Additionally, for image classification tasks, InstructBLIP (Dai et al., 2024) uses HatefulMemes (Kiela et al., 2020), a binary classification dataset for detecting hateful content in memes, to generate an instruction-response dataset. X-InstructBLIP (Panagopoulou et al., 2023) further extends this approach to audio classification by using AudioSet (Gemmeke et al., 2017). The instructions used are sentences asking for the classification of the audio, such as ”Classify the following audio”. And the response is the class of the original data.
@@ 5.1.3. SFT Data Selection
External model-based methods frequently utilize external models to evaluate the quality, diversity, and complexity of data. These models serve various functions, from scoring data to enhancing its characteristics. For instance, Du et al. (2023) leverage DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023f) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023e) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023c) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster. Zheng et al. (2024) and Sun et al. (2024) leverage LLMs to automatically select high-quality domain data, achieving superior performance. For multimodal instruction tuning datasets, MLLMs are typically used to obtain or select high-quality instruction tuning data. LLaVA1.5 (Liu et al., 2023c) pioneers the use of text-only GPT-4 to expand the COCO (Lin et al., 2014) bounding box and caption dataset into a multimodal instruction-following dataset. Following LLaVA, ShareGPT-4v (Chen et al., 2023e) utilizes GPT-4v to enhance a portion of LLaVA’s instruction tuning data. For Video Large Language Models (VideoLLMs), Liang et al. (2024) pioneered the use of CLIP score for video keyframe data selection, achieving SoTA performance.
@@
For self-instruction-based methods, the evaluation and selection of data do not require the involvement of any external models. Li et al. (2023l) proposes a self-instruction method for LLMs to autonomously identify and select challenging data. Li et al. (2023d) identifies high-quality data as those that, when included as part of a few-shot example, can enhance the model’s performance. Kung et al. (2023) proposes a novel task-level uncertainty metric that measures the sensitivity of LLMs to instruction perturbations for a task, to identify high-quality data, and combines this with the concept of active learning to iteratively select data. Liu et al. (2024c) scores data at three granular levels: token, sentence, and model. Each level’s scoring is calculated based on the previous level, with the model-level score serving as the final score.
@@ 5.2. Data-Centric Human Preference Alignment
These prompts encompass multiple tasks such as text generation, question answering, dialogue, summarization, and more. Human taggers are then asked to rank the responses to each prompt, constructing a Q&A dataset used to train the reward model (RM). Subsequently, the fine-tuning stage of the LLM is framed as a reinforcement learning (RL) problem using the trained reward model. In LLaMa2 (Li et al., 2023f), human preferences are further divided into fine-grained aspects, with two separate reward models trained to rank helpfulness and safety, respectively. Further work also aims to improve the models’ generation abilities from these aspects.
@@
To offer a comprehensive data-centric viewpoint for assessing MLLMs, this section begins with two key aspects. First, we review widely-used methods for evaluating data quality. Then, we provide an overview of MLLMs evaluation datasets, highlighting their collection, processing, and characteristics.
@@ 6.1.1. Dataset Diversity
Dataset diversity is crucial for developing robust and generalizable machine learning models. A diverse dataset enables the model to manage various scenarios and minimizes the risk of bias. Some studies (Heusel et al., 2017; Sajjadi et al., 2018) require a reference distribution or a dataset. For instance, Heusel et al. (2017) measures the Wasserstein-2 distance between two Gaussian distributions: one fitted to the embeddings of the reference sample and the other to the embeddings of the sample being evaluated for diversity. Other work use similarity scores to define diversity. For example, Shen et al. (2019); Fomicheva et al. (2020) utilize the average pairwise similarity score or its complement, the average dissimilarity. Fomicheva et al. (2020) propose a two-metric evaluation paradigm using precision and recall, with precision measuring quality and recall assessing diversity in terms of coverage of the reference distribution. The Vendi score (Friedman and Dieng, 2023; Pasarkar and Dieng, 2023; YEH et al., 2023) stands out from existing diversity evaluation metrics as a reference-free, flexible, and interpretable metric that measures internal diversity without comparing it to a reference distribution. Its reliance on a user-defined similarity function allows for broad applicability across domains, incorporating correlations between features while being computationally efficient and unsupervised, which makes it a valuable tool for diverse machine learning applications. Oppositely, Lee et al. (2023) use diversity score called Task2Vec diversity coefficient (Miranda et al., 2022) to evaluate the diversity of publicly available datasets.
@@ 6.1.2. Dataset Quality
Dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables models to learn effectively and make precise predictions. Ensuring dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enable effective learning and precise predictions. There are several approaches to evaluate data quality. TRUE (Honovich et al., 2022a) offers a thorough evaluation of factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, it highlights the robust performance of large-scale NLI and QG-QA methods across different tasks. Object Hallucination (Rohrbach et al., 2018) introduces CHAIR (Caption Hallucination Assessment with Image Relevance), which measures the proportion of words in a generated caption that correspond to objects present in the image, using ground truth sentences and object segmentations. FAITH SCORE (Jing et al., 2023) assesses the faithfulness of generated answers from large vision-language models (LVLMs). It identifies descriptive statements, extracts atomic facts, and checks their consistency with input images. This structured process ensures that generated answers align closely with the visual content, providing a comprehensive evaluation of faithfulness in LVLM outputs.
@@ 6.1.3. Dataset Similarity
In machine learning theory, performance improves when the training data distribution closely matches the testing data distribution. Therefore, it is essential to evaluate the distances of data distribution both theoretically and empirically. Common metrics for measuring distribution similarity include Euclidean distance, KL division, CORAL loss (Sun and Saenko, 2016), Wasserstein distance, and MMD distance (Jiang et al., 2022). The MAUVE scores (Pillutla et al., 2023) serve as a comparison measure between pairs of distributions, such as those encountered in generative modeling of text or images. This metric provides statistical bounds and extensive experiments to demonstrate its effectiveness.
@@ 6.2. Evaluation Datasets for MLLMs
In this section, we will briefly summarize different types of evaluation datasets according to downstream tasks.
Common evaluation tasks for MLLMs can be divided into captioning tasks, question answering tasks, perception and reasoning tasks, and other noteworthy tasks such as classification tasks. Captioning datasets (Chen et al., 2015; Venugopalan et al., 2015; Kim et al., 2019) act as benchmarks to evaluate a model’s fundamental ability to understand information from different modalities. They can be used to assess the zero-shot or few-shot learning capabilities of MLLMs in multimodality tasks, offering insights into their ability to generalize and adapt to new scenarios with limited training examples. While evaluation datasets for question answering (QA) are crucial for assessing the performance of MLLMs that handle diverse data types like text, images, and audio (Goyal et al., 2019; Xu et al., 2017; Lipping et al., 2022). These datasets require models to analyze and combine information from multiple sources to respond to relevant questions accurately. Perception and reasoning are crucial tests for assessing the general abilities of machine learning language models (MLLM). These tasks requires model to reason for complicated logic and capture structured or spatial information (Singh et al., 2019; Masry et al., 2022; Liu et al., 2023b; Mangalam et al., 2024). There are also several noteworthy tasks not mentioned above, including classification task that require model to classify examples for right labels (Kiela et al., 2020; Piczak, 2015, 2015; Wu et al., 2015; Goyal et al., 2017), and multimodal dialog task that require for model’s in-context learning abilities to answer questions during multiple rounds of conversations (Bai et al., 2023b).
@@
As elaborated in Section 6.1, despite the introduction of various data evaluation metrics, there remains a notable absence of comprehensive metrics tailored specifically for evaluating multimodal data. The assessment of multimodal data presents heightened challenges owing to its diverse array of data types, often encompassing multiple tasks and modalities (Liu et al., 2023c; Chen et al., 2024c). The quality of such data can be impacted by the individual models corresponding to each modality and their alignments. Hence, it becomes imperative to scrutinize the quality of multimodal data from multiple dimensions. In addition to relying solely on human-defined features, enhancing data quality can be achieved by leveraging statistical metrics and methodologies, such as scrutinizing distributions. The notion of domain adaptation offers theoretical underpinnings for matching distributions (Jiang et al., 2022). Consequently, assessing distributions emerges as a promising avenue, both in terms of theoretical frameworks and empirical validation.